Miller Company’s most recent contribution format income statement is shown below:

Total Per Unit
Sales (39,000 units) $273,000 $7.00
Variable expenses 156,000 4.00
Contribution margin 117,000 $3.00
Fixed expenses 46,000
Net operating income $ 71,000


Required:

Prepare a new contribution format income statement under each of the following conditions (consider each case independently): (Do not round intermediate calculations. Round your "Per unit" answers to 2 decimal places.)

   

1. The number of units sold increases by 15%.

2. The selling price decreases by $1.30 per unit, and the number of units sold increases by 19%.

3. The selling price increases by $1.30 per unit, fixed expenses increase by $9,000, and the number of units sold decreases by 4%.

4. The selling price increases by 10%, variable expenses increase by 30 cents per unit, and the number of units sold decreases by 7%.
Answer #1

Solution :-

1. The number of units sold increases by 15%. Therefore, new no. of units would be 39000*115% = 44850 Units

Total Per unit
Sales (44850 units) $313950 $7.00
Less :Variable expenses $179400 $4.00
Contribution margin $134550 $3.00
Less :Fixed expenses $46000
Net operating income $88550

2.The selling price decreases by $1.30 per unit, and the number of units sold increases by 19%. Therefore, new no. of units would be 39000*119% = 46410 Units and new selling price = $7.00 - $1.30 = $5.70

Total Per unit
Sales (46410 units) $264537 $5.70
Less :Variable expenses $185640 $4.00
Contribution margin $78897 $1.70
Less :Fixed expenses $46000
Net operating income $32897

3.The selling price increases by $1.30 per unit, fixed expenses increase by $9,000, and the number of units sold decreases by 4%. Therefore, new no. of units would be 39000*96% = 37440 Units and new selling price = $7.00 + $1.30 = $8.30 and the new fixed expenses = $46000 + $9000 = $55000.

Total Per unit
Sales (37440 units) $310752 $8.30
Less :Variable expenses $149760 $4.00
Contribution margin $160992 $4.30
Less :Fixed expenses $55000
Net operating income $105992

4. The selling price increases by 10%, variable expenses increase by 30 cents per unit, and the number of units sold decreases by 7%. Therefore, new no. of units would be 39000*93% = 36270 Units and new selling price = $7.00 *110% = $7.70 and the new variable expenses = $4 *130% =$5.20.

Total Per unit
Sales (36270 units) $279279 $7.70
Less :Variable expenses $188604 $5.20
Contribution margin $90675 $2.50
Less :Fixed expenses $46000
Net operating income $44675